OpenStudy (anonymous):

Find the slope of a line parallel to each of the following lines. A. y-x=-3 B. 5y+3x=1 C. 2x-5y+10=0

OpenStudy (anonymous):

Y=MX+B FORM?

OpenStudy (anonymous):

Yup that is the first step

OpenStudy (anonymous):

First rearrange the equation into that form

OpenStudy (anonymous):

y=x+3

OpenStudy (anonymous):

Now, do I subtract 3?

OpenStudy (anonymous):

Well we know that the slope of a parallel line will have the same slope as the original line

OpenStudy (anonymous):

so all ya gotta do is fine the slope of this line

OpenStudy (anonymous):

Now your equation is in the form y=mx+b where m=slope

OpenStudy (anonymous):

Yeah, I know

OpenStudy (anonymous):

So what is m in this equation?

OpenStudy (anonymous):

Btw you are wrong the first equation becomes y=x-3

OpenStudy (anonymous):

The m is stil the same in this case. SO in reality it wont change the answer

OpenStudy (anonymous):

So what is our m in y=x-3?

OpenStudy (anonymous):

x?

OpenStudy (anonymous):

Nopeeee y=mx+b y=1x-3 So in this case our m=1 and b=-3

OpenStudy (anonymous):

oh, so I add 3?

OpenStudy (anonymous):

can you show how u got that?

OpenStudy (anonymous):

OKKKKKKKK listen up kid

OpenStudy (anonymous):

We moved our equation into the form y=mx+b

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so we started off with y-x=-3 and then we converted it into y=x-3

OpenStudy (anonymous):

y=1x-3 is in the form y=mx+b

OpenStudy (anonymous):

Where our m=1 and b=-3

OpenStudy (anonymous):

Okk lemme give you another example Lets say we have the equation y=3x+5 The equation is in the form y=mx+b where m=3 and b=5

OpenStudy (anonymous):

Do you follow?

OpenStudy (anonymous):

Basically the m is the number in front of the x

OpenStudy (anonymous):

Yes, I get it and sorry, I was away for a couple minutes...

OpenStudy (anonymous):

so m is our slope so for our first equation the slope is 1

OpenStudy (anonymous):

ok, question 2 know

OpenStudy (anonymous):

ok first put it into y=mx+b form

OpenStudy (anonymous):

5y=3x+1

OpenStudy (anonymous):

3-3 cross out; 5-3=2

OpenStudy (anonymous):

2y=1/2

OpenStudy (anonymous):

whatttt what is 3-3 crossout???

OpenStudy (anonymous):

subtract 3 from 5 and from 3

OpenStudy (anonymous):

We are trying to isolate y Sooo we have 5y=-3x+1 Now we divide the whole equation by 5 and what do we get?

OpenStudy (anonymous):

So, I was wrong?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

I got 3/5

OpenStudy (helpblahblahblah):

me what?

OpenStudy (anonymous):

y-x=-3

OpenStudy (helpblahblahblah):

i dunno sorry

OpenStudy (anonymous):

it's ok

OpenStudy (helpblahblahblah):

okay